OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $78 for 1 adult ticket and 2 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 78 =One half x + (x - 6) What is the regular price of a child's ticket? $50 $56 $75 $81

OpenStudy (anonymous):

Just solve the equation: x/2 + x - 6 = 78 3x/2 = 78 + 6 = 84 x = 84 * 2/3 I think you can do it now.